OpenStudy (anonymous):

The first four terms of a sequence are shown below: 9, 5, 1, -3 Which of the following functions best defines this sequence?

Mehek (mehek14):

what happens to 9 that gets you to 5?

Mehek (mehek14):

so do you think this is subtraction or addition?

OpenStudy (anonymous):

(A) f(1) = 9, f(n + 1) = f(n) - 4; for n >=1 (B) f(1) = 9, f(n + 1) = f(n) + 4; for n >= 1 (C) f(1) = 9, f(n + 1) = f(n) - 5; for n >=1 (D) f(1) = 9, f(n + 1) = f(n) + 5; for n >=1 I think it's subtraction

OpenStudy (anonymous):

I also think the answer is C

Mehek (mehek14):

subtraction by 5 or by 4?

Mehek (mehek14):

ok you said C so lets check

Mehek (mehek14):

9 - 5 = 5?

OpenStudy (anonymous):

no

Mehek (mehek14):

so 9-x=5 what is x?

OpenStudy (anonymous):

I thought that f(1) is equal to 1 and up

OpenStudy (anonymous):

x = 4

Mehek (mehek14):

no they give you f(1)=9

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

So then the answer is A

Mehek (mehek14):

yup

OpenStudy (anonymous):

thank you so much!

Mehek (mehek14):

yw